a side of one of two similar triangles is four times as long as the corresponding side of the other. If the area of the larger triangle is 12 cm^2, what is the area of the smaller triangle
Answer by Cromlix(4381)

Area of larger triangle = 12 cm^2
Ratio is 1:4
Going from large area to find small area
so factor = 1/4
As you are dealing with an area,
Square factor = 1/16
12 * 1/16 = 12/16 = 0.75
Area of smaller triangle = 0.75 cm^2
Hope this helps.
:-)